# Learning to Generate Text in Arbitrary Writing Styles

### 28 Dec 2023 v1

@@ Abstract
Prior work in style-controlled text generation has focused on tasks such as emulating the style of prolific literary authors, producing formal or informal text, and the degree of toxicity of generated text. Plentiful demonstrations of these styles are available, and as a result modern language models are often able to emulate them, either via prompting or discriminative control. However, in applications such as writing assistants, it is desirable for language models to produce text in an author-specific style on the basis of a small writing sample. We find that instruction-tuned language models can struggle to reproduce author-specific style demonstrated in a prompt. Instead, we propose to guide a language model to generate text in a target style using contrastively-trained representations that capture stylometric features. A central challenge in doing so is that an author’s writing is characterized by surprising token choices under a generic language model. To reconcile this tension, we combine generative re-scoring to achieve an author-specific model, with discriminative control to ensure style consistency at the sequence-level. The combination of these approaches is found to be particularly effective at adhering to an author-specific style in a variety of conditions, including unconditional generation and style transfer, and is applicable to any underlying language model without requiring fine-tuning.
@@ 1 Introduction
We consider the problem of generating text in the style of an arbitrary author on the basis of a small writing sample, on the order of a few hundred words. Although instruction-tuned language models (LM) have demonstrated the ability to emulate a variety of writing styles via prompting Deshpande et al. (2023), particularly when a given style is well-represented in the training data Krishna et al. (2020), we find that performance is less consistent in our few-shot setting, with recent large LMs such as GPT-3.5 performing worse than previous generations Ouyang et al. (2022a). A separate challenge is that large LMs can be computationally prohibitive in certain applications, such as on-device deployment where privacy-preserving personalized generation may be needed.
@@
Prior work in controllable text generation has primarily focused on categorical target attributes such as sentiment, formality, and topic, for which a number of techniques have been proposed Prabhumoye et al. (2018); Sudhakar et al. (2019)—we discuss related work in more detail in §missing 5. However, author-specific textual styles cannot be summarized using a closed set of binary or categorical attributes, since authors may be characterized by unique combinations of stylometric features. Such features may include dialect, use of emojis, punctuation and capitalization usage, as well as less obvious features such as syntactic preferences and use of white space. Since it is difficult even for forensic linguists to characterize an author’s style, we propose to guide generation using contrastively-trained representations that extract stylistic attributes from a given writing sample as a dense vector feature. [^1] A conceptually similar approach is used in certain voice synthesis systems, in which speaker representations guide qualities of the generated speech Fang et al. (2019); Ao et al. (2021).
@@
Discriminative control methods generate text with prescribed attributes guided by a classifier evaluating the degree to which the text satisfies the target attribute, typically with a tunable hyper-parameter balancing the fluency of the generated text with control success Dathathri et al. (2019). However, human writing is characterized by “dips” into low-probability regions, unlike samples from LMs which produce likely tokens at each step Gehrmann et al. (2019). Thus, the objective of achieving fluent generations according to a generic LM will in general be in tension with the goal of matching an author’s style, which may be characterized by such unlikely token choices. To overcome this challenge, we propose a novel approach which combines a style-controlled autoregressive language model and a discriminative objective which aims to ensure stylistic consistency at the sequence level.
@@
To guide a pre-trained LM towards a target style, we generalize future discriminators Yang and Klein (2021) to regression for a target style representation. Simply put, our approach entails re-scoring the predictive distribution of an existing LM using a lightweight model that assigns higher likelihood to tokens that are predicted to better adhere to the target style vector. The resulting author-specific LM—the composition of a pre-trained model and a lightweight regressor—is then used as a fluency scorer for a discriminative model which captures stylistic consistency at the document-level. Our discriminative control framework adopts a product-of-experts energy parametrization Hinton (2002); Mireshghallah et al. (2022), which optionally enables the inclusion of meaning preservation terms in the case of style transfer (§missing 4.2).
@@
In summary, the proposed approach enables both style-controlled generation and style transfer using a pre-trained LM, without further fine-tuning. Our recipe calls for two main ingredients: a style representation (§missing 2.2) and unlabeled data to fit a lightweight re-scoring model. Since style representations are effective in various domains and unlabeled data is generally easy to come by, our approach is quite widely applicable. We conduct an extensive experimental evaluation of the proposed approach in §missing 4, finding that:
  • The proposed approach is competitive with instruction-tuned large language models, prompted to use in-context demonstrations of the target style.
@@
  • Interpolating between two target author style vectors and generating text at intermediate points yields interpretable results, considering the rate of capitalization and punctuation usage. This result suggests that our control vectors capture intuitive stylistic features and that proposed approach can then successfully reproduce those features in generated text at the expected rate.
  • Samples from the proposed approach circumvent machine generated text detectors at a higher rate, but misuse concerns can be addressed with producing more in-domain detection data.
@@
  • We show that our proposed style transfer approach can be adapted to serve as an effective author anonymization technique, defeating authorship attribution while preserving meaning.
@@ 2.1 Problem statement
We consider both sequence generation and sequence-to-sequence generation, where in both cases our objective is to produce text ($x\in\mathcal{X}$) in a target style while satisfying other criteria, such as diverse outputs in the case of language modeling and meaning preservation in the case of style transfer. We assume a few-shot setting where the target style is specified by a writing sample $y=(y_{1},y_{2},\ldots)$ exhibiting the desired stylistic attributes. In our experiments, we focus on the case where each $y_{1},y_{2},\ldots\in\mathcal{X}$ correspond to a short documents (e.g., social media comments), and we are interested in reproducing the underlying author’s specific writing style. We emphasize the difficulty of this task, stemming not only from the few-shot setting, but also the fact that stylometric features comprise a sparser signal than other more evident textual attributes like sentiment. [^2] Authorship analysis often assumes access to large corpora by the candidate authors, as in the seminal work by Mosteller and Wallace (1963). In contrast, we extract features from a relatively small number of short documents, on average comprising 68 words each.
For sequence generation, we produce text by sampling from a pre-trained LM $p$ conditioned on $y$. In the case of instruction-tuned LMs, $y$ will be paired with an appropriate prompt to elicit the desired output; we discuss prompting strategies in more detail in §missing 4.1.2. In sequence-to-sequence generation, we are additionally given initial text $x^{(0)}$ that we wish to revise to be closer to the target style $y$, while keeping other properties of $x^{(0)}$ constant, such as preserving meaning. Rather than condition on $y$ directly as done in the prompting approach, we propose instead using a discriminative feature extractor $f$ to capture stylistic properties of $y$, which are then used to guide generation. This is a distinguishing characteristic of our approach, since much prior work in controllable text generation has focused on classifiers (e.g., sentiment polarity) and prompting strategies to guide generation. We discuss the feature extractor in more detail next.
@@ 2.2 Author style representations
As previously mentioned, author-specific style is difficult to characterize even for forensic linguists, which poses challenges both for control and for evaluation. However, recent work has leveraged the availability of large corpora of writings by anonymous authors to learn stylistic representations. Although lacking in interpretability, such representations have been found to be effective at discriminating between authors, a task which requires characterizing writing style Wang et al. (2023). In this work, we consider two different representations, both trained for surrogate tasks of authorship prediction. The representations aim to capture stylistic features of authorship, which is achieved by training on corpora with sufficient topic diversity, or by specifically mining examples of the same author writing about different topics (and different authors writing about the same topic). To avoid “gaming” our evaluation metrics, we employ different models for control and for evaluation.
@@
To guide generation, we adapt the model proposed by Rivera-Soto et al. (2021). Specifically, we estimate a representation $f$ on the basis of a large collection of anonymous writing samples. Our training dataset consists of one million Reddit users, each contributing at least 100 comments Khan et al. (2021). The unique account labels enable supervised contrastive training, encouraging features $f(x)$ and $f(x^{\prime})$ to be similar when $x$ and $x^{\prime}$ have the same author. [^3] We use code provided by Rivera-Soto et al. (2021) at https://github.com/LLNL/LUAR.
@@
We use two models for evaluation which are both available publicly as pre-trained checkpoints (§missing 4.1.1). The first is a further instance based on the recipe from Rivera-Soto et al. (2021) trained on a larger corpus of 5 million authors, resulting in a more capable model than the one we use to guide generation. We also use a model proposed by Wegmann et al. (2022), which is trained on different data and using a different training objective. Although substantially less effective in author verification settings, this model uses topic labels to attempt to produce representations that are less sensitive to topical similarity.
@@ 3 Guiding generations towards a target style representation
In this section, we describe our approach, which aims to generate text $x$ satisfying various soft constraints, the most important of which is adherence to the style demonstrated in the few-shot example $y$. To reconcile the tension between fluency and author-specific style, we first show how to use a regression model to guide an LM to produce text for which $f(x)$ is close to $f(y)$ in expectation. Next, we show how the resulting author-specific LM can be incorporated in an energy-based model (EBM) using a product-of-experts, which confers two advantages. First, the EBM is a non-autoregressive model which performs inference at the sequence-level; therefore, the distance between $f(x^{\prime})$ and $f(y)$ can be directly evaluated to score candidate generations $x^{\prime}$. Second, this framework makes it straighforward to introduce further experts to satisfy arbitrary additional preferences, such as meaning preservation in the case of style transfer.
3.1 Future regressors
Figure 1: Style performance for an increasing number of examples of a target style. We find that more examples result in better representations, which in turn improve decoding quality. Our future regressor and EBM approaches significantly outperform much larger models using prompting strategies.
@@ 3.1 Future regressors
This approach affords a natural extension to continuous control by fitting a future regressor $p(f(x)=\mathbf{c}\mid x_{1:i})$, where $f(x)$ is evaluated on the sequence and the model is conditioned on all prefixes of the same sequence, and therefore learns to predict the probability that a given prefix $x_{1:i}$ will adhere to the target style in the future. To do so, we stipulate that control vectors c are distributed according to a multivariate Normal density,
$\textbf{c}\mid x_{1:i}\sim\text{Normal}(\mu,\Sigma)$
and parameterize $\mu$ and $\Sigma$ using neural networks with input $x_{1:i}$ and where $\Sigma$ constrained to be a valid covariance matrix. The simplest approach to enforce this constraint is to assume all off-diagonal entries of $\Sigma$ are zero with positive values along the diagonal, which we achieve with a softplus transformation of the unconstrained final activations of the neural network. Irrespective of the output parameterization, we employ a shared network $g_{\theta}$ for both $\mu$ and $\Sigma$; specifically, for the diagonal covariance, we define:
z $=g_{\theta}(x_{1:i})$
@@
The parameters $\Theta=(\theta,\phi,\eta)$ are optimized on the basis of a corpus consisting of text paired with corresponding control vectors $\{(\textbf{c},x)\}_{i=1}^{N}$. In general, and in the experiments reported in this paper, $g_{\theta}$ will have many fewer parameters than the LM being guided, in which case evaluating $p(\mathbf{c}\mid x_{1:i-1})$ during generation introduces a relatively small additional computational burden. Implementation details for this architecture are outlined in §missing 4.
The diagonal covariance matrix assumption stipulates that each component of a control vector is independent. While previous work in contrastive learning has found that explicitly enforcing decorrelation to be necessary for such an assumption to be effective Tao et al. (2021), we find that the control vectors we consider already satisfy this condition quite well. [^4] In fact, we trained a model using the decorrelation objective and found the associated control vectors yielded no noticeable improvement in downstream decoding.
Optimization For each training instance $(\textbf{c},x)$, we create the augmented set consisting of all prefixes $(\textbf{c},x_{1:1})$, $(\textbf{c},x_{1:2})$, $\ldots$, $(\textbf{c},x_{1:n})$. Note that the target c is the same for each prefix, since the regressor is predicting whether the target control vector will be true for the full $x$ on the basis of the supplied prefix. The parameters $\Theta$ of the regression model are optimized to maximize the log-likelihood of the observed control vectors. We found it effective to initialize $f_{\theta}$ using the same model that extracted the reference control vectors  [^5] We found that using this initialization resulted in a 1.4% performance improvement over a random initialization., before fine-tuning $\Theta$ on the augmented data.
3.2 An energy-based model for non-autoregressive generation
The proposed future regressor can be combined with any underlying LM to produce samples $x$ with stylistic features $f(x)$ close to the target $f(y)$ in expectation. However, autoregressive generation incrementally constructs the sample $x$, and therefore cannot directly use the feature-space distance between $f(x)$—based on the complete sample $x$—and the target $f(y)$, to guide generation. Additionally, to support tasks such as style transfer (§missing 4.2), it is necessary to impose additional arbitrary constraints on generation such as meaning preservation, which is not easy to do in a Bayesian formulation Yang and Klein (2021).
@@ 3.2 An energy-based model for non-autoregressive generation
with experts $E_{i}$ corresponding to soft constraints; this model assigns higher probability to sequences $x$ which simultaneously satisfy all constraints. Note that evaluating the above probability requires an intractable sum over all possible sequences $x$; therefore, we resort to approximate inference (§missing 3.3). We consider two settings in our experiments: style-controlled generation and style transfer. In both settings, we have found it effective to tune the weights $\alpha$ using validation data, although we note that maximum-likelihood estimation may also be used by approximating the gradient, to avoid any manual tuning.
@@
Here we use only two experts. $E_{1}$ is an author-specific LM as described in the previous section, which evaluates the negative log-probability of $x$ under the author-adapted LM. $E_{2}$ is an expert measuring sequence-level style similarity. Specifically, $E_{2}$ computes the distance between the style vector of $x$ and a target style control vector $f(y)$ via the negative angular similarity,
$-\left(1-\text{arccos}\left(\frac{f(x)\cdot f(y)}{\|f(x)\|\|f(y)\|}/\pi\right)\right).$

### 04 Mar 2024 v2

@@ Abstract
Prior work in style-controlled text generation has focused on tasks such as emulating the style of prolific literary authors, producing formal or informal text, and mitigating toxicity of generated text. Plentiful demonstrations of these styles are available, and as a result modern language models are often able to emulate them, either via prompting or discriminative control. However, in applications such as writing assistants, it is desirable for language models to produce text in an author-specific style on the basis of a potentially small writing sample. For example, someone writing in a particular dialect may prefer writing suggestions that retain the same dialect. We find that instruction-tuned language models can struggle to reproduce author-specific style demonstrated in a prompt. Instead, we propose to guide a language model to generate text in a target style using contrastively-trained representations that capture stylometric features. Our approach (StyleMC) combines an author-adapted language model with sequence-level inference to improve stylistic consistency, and is found to be effective in a variety of conditions, including unconditional generation and style transfer. Additionally, we find that the proposed approach can serve as an effective anonymization method, by editing a document to mask authorship while preserving the original meaning.
@@ 1 Introduction
We consider the problem of generating text in the style of an arbitrary author on the basis of a small writing sample, on the order of a few hundred words. Although instruction-tuned language models (LM) have demonstrated the ability to emulate a variety of writing styles via prompting Deshpande et al. (2023), particularly when a given style is well-represented in the training data Krishna et al. (2020), we find that performance is less consistent in our few-shot setting, with recent large LMs such as GPT-3.5 performing worse than its previous generations. A separate challenge is that large LMs can be computationally prohibitive in certain applications, such as on-device deployment where privacy-preserving personalized generation may be needed.
Figure 1: An overview of StyleMC for style transfer. Using MCMC, we generate text based on a few samples of the target style, an author-specific fluency model (§3.2), and the original content. Our approach reproduces salient features of an author’s style while preserving meaning. In the (real) example above, accepted samples exhibit characteristics of British English (yellow), matching the characteristics of the target style.
Prior work in controllable text generation has primarily focused on categorical target attributes such as sentiment, formality, and topic, for which a number of techniques have been proposed Prabhumoye et al. (2018); Sudhakar et al. (2019); Yang and Klein (2021)—we discuss related work in more detail in §missing 7.
@@
However, author-specific textual styles cannot be summarized using a closed set of binary or categorical attributes, since authors may be characterized by unique combinations of stylometric features. Such features may include dialect, use of emojis, punctuation and capitalization usage, as well as less obvious features such as syntactic preferences and use of white space. Since it is difficult even for forensic linguists to characterize an author’s style, we propose to guide generation using contrastively-trained representations that extract stylistic attributes from a given writing sample as a dense vector feature. [^1] A conceptually similar approach is used in certain voice synthesis systems, in which speaker representations guide qualities of the generated speech Fang et al. (2019); Ao et al. (2021).
@@
Discriminative control methods generate text with prescribed attributes guided by a classifier evaluating the degree to which the text satisfies the target attribute, typically with a tunable hyper-parameter balancing the fluency of the generated text with control success Dathathri et al. (2019). However, human writing is characterized by “dips” into low-probability regions, unlike samples from LMs which produce likely tokens at each step Gehrmann et al. (2019). Thus, the objective of achieving fluent generations according to a generic LM will in general be in tension with the goal of matching an author’s style, which may be characterized by such unlikely token choices. To overcome this challenge, we propose StyleMC, a novel approach which combines a style-controlled autoregressive language model and a discriminative objective which aims to ensure stylistic consistency at the sequence level.
@@
To guide a pre-trained LM towards a target style, we generalize future discriminators Yang and Klein (2021) to regression for a target style representation. Simply put, our approach involves re-scoring the predictive distribution of an existing LM using a lightweight model that assigns higher likelihood to tokens that are predicted to better adhere to the target style vector. The resulting author-specific LM—the composition of a pre-trained model and a lightweight regressor—is then used as a fluency scorer for a discriminative model which captures stylistic consistency at the document-level. Our discriminative control framework adopts a product-of-experts energy parametrization Hinton (2002); Mireshghallah et al. (2022), which enables style transfer through the inclusion of meaning preservation terms (§missing 5).
@@
In summary, StyleMC enables both style-controlled generation and style transfer using a pre-trained LM, without further fine-tuning. Our recipe calls for two main ingredients: a style representation (§missing 2.2) and unlabeled data to fit a lightweight re-scoring model. Since style representations are effective in various domains and unlabeled data is generally easy to come by, our approach is quite widely applicable. We conduct an extensive experimental evaluation of the proposed approach in §missing 4, §missing 5, and §missing 6, finding that:
  • StyleMC is proficient in style generation tasks (including style transfer), outperforming large language models such as GPT-4 that were prompted to use in-context demonstrations of the target style.
@@
  • Interpolating between two style vectors and generating text at intermediate points using our approach yields interpretable results, considering the rate of capitalization and punctuation usage. This result suggests that (a) our control vectors capture intuitive stylistic features; and (b) that the proposed approach can successfully reproduce those features in generated text at the expected rate.
@@
  • Our proposed style transfer approach can be adapted to serve as an effective author anonymization technique, defeating authorship attribution while preserving meaning.
  • In a zero-shot setting, samples from the proposed approach are harder to detect as being machine-generated than other LLMs, which we attribute to a greater ability to mimic human writing style.
Reproducibility
We release reference model implementations, checkpoints, datasets, and experiment scripts. Commodity hardware (e.g., a single GPU) is sufficient to reproduce most of our results.
@@ 2.1 Problem statement
We consider both sequence generation (§missing 3.1) and sequence-to-sequence generation (§missing 3.2), where in both cases our objective is to produce text ($x$) in a target style while satisfying other criteria, such as diverse outputs in the case of language modeling and meaning preservation in the case of style transfer. We assume a few-shot setting where the target style is specified by a writing sample $y=(y_{1},y_{2},\ldots)$ exhibiting the desired stylistic attributes. In our experiments, we focus on the case where each $y_{1},y_{2},\ldots$ correspond to a short documents (e.g., social media comments), and we are interested in reproducing the underlying author’s specific writing style. We emphasize the difficulty of this task, stemming not only from the few-shot setting, but also the fact that stylometric features comprise a sparser signal than other more evident textual attributes like sentiment. [^2] Statistical authorship analysis often assumes access to large corpora by the candidate authors, as in the seminal work by Mosteller and Wallace (1963). In contrast, we extract features from a relatively small number of short documents, on average comprising 68 words each.
For sequence generation, we produce text by sampling from a pre-trained LM $p$ conditioned on $y$. In the case of instruction-tuned LMs, $y$ will be paired with an appropriate prompt to elicit the desired output; we discuss prompting strategies in more detail in §missing 4.2. In sequence-to-sequence generation, we are additionally given initial text $x^{(0)}$ that we wish to revise to be closer to the target style $y$, while keeping other properties of $x^{(0)}$ constant, such as semantic meaning. Rather than condition on $y$ directly as done in the prompting approach, we propose instead using a discriminative feature extractor $f$ to capture stylistic properties of $y$, which are used to guide generation. This is a distinguishing characteristic of our approach, since much prior work in controllable text generation has focused on classifiers (e.g., sentiment polarity) and prompting strategies to guide generation. We discuss the feature extractor in more detail next.
@@ 2.2 Author style representations
As previously mentioned, author-specific style is difficult to characterize even for forensic linguists, which poses challenges both for control and for evaluation. However, recent work has leveraged the availability of large corpora of writings by anonymous authors to learn stylistic representations. Such representations have been found to be effective at discriminating between authors by characterizing writing style Wang et al. (2023). In this work, we consider two different representations, both trained for surrogate tasks of authorship prediction.
@@
To guide generation, we adapt the model proposed by Rivera-Soto et al. (2021). Specifically, we estimate a representation $f$ on the basis of a large collection of anonymous writing samples. Our training dataset consists of one million Reddit users, each contributing at least 100 comments Baumgartner et al. (2020); Khan et al. (2021). The unique account labels enable supervised contrastive training, encouraging features $f(x)$ and $f(x^{\prime})$ to be similar when $x$ and $x^{\prime}$ have the same author. [^3] We use code provided by Rivera-Soto et al. (2021) at https://github.com/LLNL/LUAR.
@@
We use two models for evaluation which are both available publicly as pre-trained checkpoints (§missing 4.1). The first is a further instance based on the recipe from Rivera-Soto et al. (2021) trained on a larger corpus of 5 million authors, resulting in a more capable model than the one we use to guide generation. We also use a model proposed by Wegmann et al. (2022), which is trained on different data using topic labels to attempt to produce representations that are less sensitive to topical similarity.
@@ 3 Guiding generations towards a target style representation
In this section, we describe StyleMC, which aims to generate text $x$ satisfying various soft constraints, the most important of which is adherence to the style demonstrated in the few-shot example $y$. To reconcile the tension between fluency and author-specific style, we first show how to use a regression model to guide an LM to produce text for which $f(x)$ is close to $f(y)$ in expectation. Next, we show how the resulting author-specific LM can be incorporated in an energy-based model (EBM) using a product-of-experts, which confers two advantages. First, the EBM is a non-autoregressive model which performs inference at the sequence-level; therefore, the distance between $f(x^{\prime})$ and $f(y)$ can be directly evaluated to score candidate generations $x^{\prime}$. Second, this framework makes it straightforward to introduce further experts to satisfy arbitrary additional preferences, such as meaning preservation in the case of style transfer.
3.1 Few-shot language model adaptation
@@
This approach affords a natural extension to continuous control by fitting a future regressor $p(f(x)=\mathbf{c}\mid x_{1:i})$, where $f(x)$ is evaluated on the sequence and the model is conditioned on all prefixes of the same sequence, and therefore learns to predict the probability that a given prefix $x_{1:i}$ will adhere to the target style in the future. To do so, we stipulate that control vectors c are distributed according to a multivariate Normal density, and parameterize $\mu$ and $\Sigma$ using neural networks with input $x_{1:i}$ and where $\Sigma$ constrained to be a diagonal convariance. Specifically, we employ a shared network $\textbf{z}=g_{\theta}(x_{1:i})$ for both $\mu$ and $\Sigma$:
@@
The parameters $\Theta=(\theta,\phi,\eta)$ are optimized on the basis of a corpus consisting of text paired with corresponding control vectors $\{(\textbf{c},x)\}_{i=1}^{N}$. In general, and in the experiments reported in this paper, $g_{\theta}$ will have many fewer parameters than the LM being guided, in which case evaluating $p(\mathbf{c}\mid x_{1:i-1})$ during generation introduces a relatively small additional computational burden. Implementation details for this architecture are outlined in §missing 4.2.
The diagonal covariance matrix implies that each component of the control vector is independent. While previous work in contrastive learning has found that explicitly enforcing decorrelation to be necessary for such an assumption to be effective Tao et al. (2021), we find that the control vectors we consider already satisfy this condition quite well. [^4] In fact, we trained a model using the decorrelation objective and found the associated control vectors yielded no noticeable improvement in downstream decoding.
Optimization For each training instance $(\textbf{c},x)$, we create the augmented set consisting of all prefixes $(\textbf{c},x_{1:1})$, $(\textbf{c},x_{1:2})$, $\ldots$, $(\textbf{c},x_{1:n})$. Note that the target c is the same for each prefix, since the regressor is predicting whether the target control vector will be true for the full $x$ on the basis of the supplied prefix. The parameters $\Theta$ of the regression model are optimized to maximize the log-likelihood of the observed control vectors. We found it effective to initialize $g_{\theta}$ using the same model that extracted the reference control vectors  [^5] We found that using this initialization resulted in a 1.4% performance improvement over a random initialization., before fine-tuning $\Theta$ on the augmented data.
3.2 A unified sequence-level model for style control and style transfer
The proposed future regressor can be combined with any autoregressive LM to produce samples $x$ with stylistic features $f(x)$ close to the target $f(y)$ in expectation. However, autoregressive generation incrementally constructs the sample $x$, and therefore cannot directly use the feature-space distance between $f(x)$—based on the complete sample $x$—and the target $f(y)$, to guide generation. Additionally, to support tasks such as style transfer (§missing 5), it is necessary to impose additional constraints on generation such as meaning preservation.
@@ 3.2 A unified sequence-level model for style control and style transfer
with experts $E_{i}$ corresponding to soft constraints; this model assigns higher probability to sequences $x$ which simultaneously satisfy all constraints. Since evaluating the above probability requires an intractable sum over all possible sequences $x$, we resort to approximate inference (§missing 3.3). We consider two settings in our experiments: style-controlled generation and style transfer. In both settings, we have found it straighforward to tune the weights $\mathbf{\alpha}$ using validation data, although we note that maximum-likelihood estimation could be used instead to avoid any manual tuning.
@@
Here we use only two experts. $E_{1}$ is an author-specific LM as described in the previous section, which evaluates the negative log-probability of $x$ under the author-adapted LM. In our implementation, we average the log-probabilities for each token rather than summing them to ensure sequence length does not skew energy scores. $E_{2}$ is an expert measuring sequence-level style similarity. Specifically, $E_{2}$ computes the distance between the style vector of $x$ and a target style control vector $f(y)$ via the negative angular similarity.To avoid noisy estimates for $f(x)$ when dealing with short text samples, we consider a candidates post within the context of other samples from the same author. Specifically, when revising a text sample $x_{i}$ from the writing sample $x=(x_{1},x_{2}...)$, our EBM computes $E_{2}(x,y)$ as opposed to $E_{2}(x_{i},y)$.
Controlled text revision In the style transfer task, we additionally condition generation on an initial state $x^{(0)}$, and the objective is to modify $x_{0}$ to adhere to the style of $y$ while preserving the original meaning of $x^{(0)}$. To do so, we employ $E_{1}$ and $E_{2}$ as before, but introduce further experts that are functions of $x$ and $x^{(0)}$ and measure meaning preservation. We note that various options are possible for this purpose and our specific choices may not be optimal in all cases. In our experiments, we refer to $E_{3}$ as the measure of semantic similarity. To ensure that $x$ makes minimal revisions to $x^{(0)}$, we additionally add $E_{4}$ defined as the Hamming distance between $x$ and $x^{(0)}$, which was also employed by Mireshghallah et al. (2022).
3.3 Inference
We frame the generation problem as finding an output $x$ which minimizes energy defined by Equation (missing) 1. Although this problem is intractable, the Metropolis-Hastings (MH) algorithm can be used to obtain an approximate sample from the desired distribution. Differing from Goyal et al. (2021), we use T5 Raffel et al. (2019), an encoder-decoder model trained with an in-filling objective, to obtain proposals for the sampling scheme. [^6] We also experimented with a masked language model and found that the proposals were significantly lower quality than those produced by T5. At each step in the procedure, a fixed number of tokens are masked which T5 in-fills with a variable number of tokens—possibly fewer than were masked—to generate a candidate state. In Appendix B we vary the proposal model and masking scheme, finding that masking two tokens at a time, and infilling with T5-3B produced the best results.
In general, the state of the sampler may consist of more than one document, which together must adhere to the target style. At each step, we sample one of the documents $i$ for a MH update uniformly at random, and make proposals according to the proposal model conditioned on $x_{i}$, but evaluate $E_{2}(x,y)$ based on the entire sampler state. Thus, the energy function captures similarities on the entire state as opposed to a single document. We also run the sampler for a fixed number of steps (80 times the length of the sequence). For decoding, we record the intermediate states at each iteration, compute the energy for each state, and select the $argmax$ as our final output.
For style transfer (§missing 5), we experiment with restrictions on our proposal distribution based on part-of-speech tags. We hypothesize that nouns most often do not characterize writing style and so can be fixed during our inference procedure without impacting the ability to control style, with the potential benefit of helping with mixing and meaning preservation. We adapt a part-of-speech tagger proposed by  Sajjad et al. (2022) and tag teach token in our initial sequence. While sampling an edit point, we do not allow the model to sample a token which is tagged as any form of noun. The improvement from disallowing any edits to nouns can be observed in our style transfer results (Table 2).
4 Style Control Experiments
4.1 Metrics
Discriminating between fine-trained styles (i.e. on a per author basis) is a challenging task for human evaluators. Therefore, our evaluation of control success relies on automatic metrics. To avoid concerns about gaming certain metrics, we include multiple automatic metrics for each text attribute that is measured. We measure the overall quality of the generated text through fluency in addition to particular features (e.g. semantic meaning, style consistency) in generated text Celikyilmaz et al. (2020). We also consider further downstream tasks to evaluate the quality of style altered text, like author detection (§missing 6.1) and LM generated text detection (§missing 6.2).
Style similarity
To measure how well generated text matches a target style, we adapt previous work (discussed in §missing 2) as automatic evaluation tools. We consider “Universal Author Representations” (UAR) Rivera-Soto et al. (2021) and “Content Independent Style Representations” (CISR) Wegmann et al. (2022) [^7] Checkpoints: https://huggingface.co/rrivera1849/LUAR-CRUD and https://huggingface.co/AnnaWegmann/Style-Embedding.. These pre-trained embeddings measure style overlap between generated and reference text samples. We report cosine similarity between reference and generated text embeddings.
Fluency
Beyond satisfying style-specific constraints, generated text should remain fluent. We emphasize however that fluency is, to some extent, at odds with the goal of introducing author-specific style. [^8] The human reference data in Table 1 has an average perplexity of 205.04 measured by Mistral-7B. That is, if the target style samples have a high perplexity under an LM, it is reasonable to expect a well formed style controlled generation to also have a high perplexity. We directly measure and report the percent difference between generated and reference fluencies throughout our results, where a smaller difference is better. We use Mistral-7B to measure fluencies of all generated text Jiang et al. (2023).
4.2 Experimental Setup
Figure 2: Style performance for an increasing number of examples of a target style. We find that more examples result in better representations, which in turn improve decoding quality. Our proposed approach, including an ablation without sequence-level decoding (- EBM), significantly outperforms much larger models using prompting strategies.
Evaluation datasets
We evaluate the effectiveness of our decoding strategy on 800 authors contributing to four unique Reddit subreddits. We consider Reddit data collected through the Pushshift API Baumgartner et al. (2020) and compile a test split for each of our 4 subreddits: /r/wsb, /r/AskHistorians, /r/news, and /r/australia; performance on all four results are reported together in Table 1. These subreddits are selected for their unique and distinctive styles. We use an additional split from /r/wsb to validate our methods (i.e. to select optimal decoding hyperparameters). For each author, we compile $N$ text samples as the source of style evidence. In all experiments except for Figure 2, we set $N=16$ to balance the quality of resulting style representations and cost of compiling data following previous work Andrews and Bishop (2019). Figure 2 illustrates part of this trade-off, with improved style controlled generation for larger $N$ values.
Language models
Our proposed methods operate on a frozen underlying LM; we do not perform any fine-tuning. Across all experiments we use variants of OPT and MPT-7B to generate text Zhang et al. (2022); Team (2023). [^9] A single NVIDIA V100 GPU was sufficient to store both the control model and language model, except for 7B parameter experiments which required a second V100 GPU. For EBMs, we use T5-3B as our proposal distribution.
Future Regressors
We train the forward looking regressor using a single V100 GPU, a batch size of 64, and a learning rate of $1e^{-6}$ for 100k steps. In our experiments we apply this re-scoring procedure to OPT-1.3B to capture author specific fluencies, and we evenly weight likelihoods from OPT and the forward regressor to compute re-scored sequence level likelihoods.
$\uparrow$ Success